A piece of wire, 360 cm long is used to make the twelve edges of a rectangular box in which the length is twice the breadth.
Denoting the breadth of the box by x cm, show that the volume of the box, V cm^3 is given by V=180x^2-6x^3.

Answer by advanced_Learner(501)

let x be width then lenght is 2x
area is A=x*2x=2x^2
lenght is made up of 4 sides so 8x
widht is made of 4 sides so 4x
total is 8x+4x =12x while 360-12x is remainng height
(360-12x)/(4)=90-3x is the height
volume=area*height=(2x^2)(90-3x)
180x^2-6x^3.